OpenStudy (anonymous):

will give medal for alg 2 help Jessica is saving money for a new sailboat. She has already saved $1,000 and plans to save $150 per week. This situation can be represented by the function f(x) = 1,000 + 150x, where x is the number of weeks. How much money has Jessica saved after 12 weeks? $1,000 $1,800 (i think its this one) $2,800 $3,000

OpenStudy (anonymous):

How did you get $1800?

OpenStudy (anonymous):

Did you plug in 12 into f(x) to evaluate?

OpenStudy (anonymous):

ok im a little confused

OpenStudy (anonymous):

Well, the function you are given: f(x) gives you the total money saved after x number of weeks right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Evaluate the expression: \[f(12) = $1000 + $150(12)\]

OpenStudy (anonymous):

f(12)=1300

OpenStudy (anonymous):

\[1000 + 150 \times 12 = 1000 + 1800 = 2800\]

OpenStudy (anonymous):

Please check your calculations.

OpenStudy (anonymous):

oh oops i put 2 instead of 12